# The Cohen–Lenstra moments over function fields via the stable homology of non-splitting Hurwitz spaces

### v2

@@ 4.5. Showing the map is an equivalence
The proof we give here is modeled off of the argument in [RW20, Corollary 5.4]. We note that the argument given there has a small gap, where it is claimed on the first line of [RW20, p. 15] that $B\operatorname{Hur}^{c}_{G}\to B\operatorname{Conf}$ being a rational homology equivalence implies $\Omega B\operatorname{Hur}^{c}_{G}\to\Omega B\operatorname{Conf}$ is a rational homology equivalence on each path component. In general this does not hold, but it does hold when the spaces are weakly simple (see Section 4.5 below), as was very helpfully pointed out to us by Oscar Randal-Williams. This observation allows us to fix the small gap.
@@
This lemma follows as in the proof of [RW20, Corollary 5.4]. Using Section A.4 for $X=c$, $M=N=*_{+}$, we obtain an explicit topological model for $B\operatorname{Hur}^{c}$. By filtering this space by the number of labels on the interior, we get a filtration such that the associated cohomology spectral sequence is the cochain complex computing rack cohomology of $c$, see [EG03]. It follows from [EG03, Theorem 4.2] that the map $c\to*$ induces an isomorphism on rack cohomology with $\mathbb{Z}[\frac{1}{G^{0}_{c}}]$-coefficients, which completes the proof. ∎
Proof of Section 4.5.
@@
We next wish to reduce to showing $B\operatorname{Hur}^{c}\to B\operatorname{Conf}^{c}$ is a rational homology equivalence. It is shown in [FRS07, Proposition 5.2] that for any rack $X$, $BX$ is weakly simple. By Section 4.5, $B\operatorname{Conf}^{c}$ is also weakly simple, since the action of $\pi_{1}$ on higher homotopy groups factors through $\pi_{1}B\operatorname{Conf}$ acting on the higher homotopy groups of $B\operatorname{Conf}$, and $B\operatorname{Conf}$ is the classifying space of a rack.
Applying Section 4.5 to the map $B\operatorname{Hur}^{c}\to B\operatorname{Conf}^{c}$, we learn that if it is a a rational homology equivalence, then $\Omega B\operatorname{Hur}^{c}\to\Omega B\operatorname{Conf}^{c}$ is also a rational homology equivalence. Here, we are also using that the map $\pi_{1}B\operatorname{Hur}^{c}\to\pi_{1}B\operatorname{Conf}^{c}$ is an equivalence by construction of $\operatorname{Conf}^{c}$.
We next claim that $B\operatorname{Conf}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology. The map $\Omega B\pi_{0}\operatorname{Hur}^{c}\to\Omega B\pi_{0}\operatorname{Conf}$ is surjective with kernel a group we call $G$. Here, $G$ is finite because of the fact that $|\pi_{0}\operatorname{Conf}^{c}_{n}|=|\pi_{0}\operatorname{Hur}^{c}_{n}|$ stabilizes as $n$ grows, which is a consequence of § 4.4 and the fact that a filtered colimit of finite sets along surjections is finite. It follows from this and the fact that $\psi$ from (4.3) is an equivalence that the fiber of the map $B\operatorname{Conf}^{c}\to B\operatorname{Conf}$ is $BG$. We then obtain our claim that $B\operatorname{Conf}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology using the Serre spectral sequence for this fiber sequence, since the rational homology of $G$ is just $\mathbb{Q}$ in degree 00 with trivial action of $B\operatorname{Conf}$.
It remains then to show that the map $B\operatorname{Hur}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology, but this follows from applying Section 4.5, since $c/c=*$ since $c$ is connected, and $\operatorname{Hur}^{*}$ is $\operatorname{Conf}$. ∎
@@
For $c^{\prime}\subset c$ a subrack, there is a map of $\mathbb{E}_{1}$-algebras $\overline{i}_{c^{\prime}}^{c}:\operatorname{Conf}^{c^{\prime}}\to\operatorname{Conf}^{c}$ that is an inclusion of components. Similarly, there is a map $\tilde{i}_{c^{\prime}}^{c}:\operatorname{Hur}^{c^{\prime}}\to\operatorname{Hur}^{c}$. We have a section of the induced map $\operatorname{Conf}^{c^{\prime}}_{+}\to\operatorname{Conf}^{c}_{+}$ of pointed $\mathbb{E}_{1}$-algebras $\overline{r}_{c}^{c^{\prime}}:\operatorname{Conf}^{c}_{+}\to\operatorname{Conf}^{c^{\prime}}_{+}$ obtained by sending all components not in image of $\operatorname{Conf}^{c}$ to the base point. Similarly there is a map $\widetilde{r}_{c}^{c^{\prime}}:\operatorname{Hur}^{c}_{+}\to\operatorname{Hur}^{c^{\prime}}_{+}$ obtained the same way. Let $r_{c}^{c^{\prime}}$ denote the associated $\mathbb{E}_{1}$-algebra map $A\to C_{*}(\operatorname{Hur}^{c^{\prime}})$ obtained by taking reduced chains. In particular, this map of algebras sends all generators $\alpha_{x}$, for $x\in c-c^{\prime}$, to 00.
@@
We first note that since $\alpha_{c-c^{\prime}}$ acts by 00 on $C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$, completion doesn’t change it, so the map $A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}\to C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$ in the proposition statement is obtained by taking $r_{c}^{c^{\prime}}$, localizing at $[\alpha_{c^{\prime}}^{-1}]$ and completing at $\alpha_{c-c^{\prime}}$.
@@
${A}$${C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]}$${A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}.}$$\scriptstyle{h}$
@@
Note that all of the objects in the triangle are connective. By § 3.3, $\pi_{0}A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ is the quotient of $\pi_{0}A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ by the two sided ideal generated by $\alpha_{c-c^{\prime}}$. This ideal is exactly the classes coming from components not in $\operatorname{Hur}^{c^{\prime}}$. Thus the map $h$ is an equivalence on $\pi_{0}$. Moreover, $A\to A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ is a hom. epi. by § 4.4 and § 3.3.
@@
is an ind-weak homology equivalence (see § A.4) as $\epsilon$ approaches 00 where $10<\epsilon<$ 1.
We now recall the description of these spaces, which follows from § A.4. First, ${Q}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$ is the space whose points consist of tuples $(m,n,x,l:x\to c)$ with the following properties: $m\in X_{+},n\in X_{+}$, $x$ is a finite subset $x$ of $[0,1]\times[\epsilon,1-\epsilon]$ such that if $\pi_{2}$ is the projection onto the second coordinate, then $\pi_{2}$ is injective on $x$, and the points in $\pi_{2}(x)$ are distance at least $\epsilon$ from each other, and $l:x\to c$ is a map of sets, which we colloquially refer to as a labeling function.
@@
Suppose $y$ as above lies in $F_{j}Q_{\epsilon}^{c,c^{\prime}}$. Define $v$ to be the largest number so that $l(p_{v})\notin N_{c}(c^{\prime})$. By definition of the filtration $F_{j}Q_{\epsilon}^{c,c^{\prime}}$, we have $v\geq k-j$. Let $\epsilon^{\prime}:=\frac{\epsilon}{2}$. We now construct a continuous homotopy $H:F_{j}Q^{c,c^{\prime}}_{\epsilon}\times I\to G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$ as follows: If either $m,n$ is the base point, the point $y$ is equivalent to the base point under relation $(iii)$, and we choose the constant homotopy at the base point. So now assume this is not the case. If $v>k-j$, then the point is in $F_{j-1}Q_{\epsilon}^{c,c^{\prime}}$, so goes to the base point in the associated graded $G_{j}Q_{\epsilon}^{c,c^{\prime}}$. Thus, we may choose the constant homotopy on such points as well. So now we may assume as well that $v=k-j$.
@@
It only remains to consider then the case that the point $p_{r}$ being removed in $(i)$ has $l(p_{r})\in c^{\prime}$. Note that we may assume $r\neq v$ because $l(p_{v})\notin c^{\prime}$. Because we are working on the subquotient $G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$, if we apply $(i)$ for some point $p_{r}$ with $r>v$, the point becomes identified with the base point, and so $H$ is the constant homotopy on the base point. (This is the reason for introducing the filtration $F_{j}$.) Hence, it only remains to deal with the case that $r<v$ at a point $y$ presented as $(m,n,x,l:x\to c)$, as in ()($\dagger$)( † ) above.
@@
First, recall the defining relation of Hurwitz space from § 4.1 implies that the positive generator of the braid group identifies $[g]\cdot[h]=[h]\cdot[h\triangleright g]$. Therefore, taking $h=\xi(l(p_{v}))$, and using that $[\xi(l(p_{v}))]$ acts invertibly on $M$, the action on $M$ factors through a group where the relation $[\xi(l(p_{v}))]^{-1}\cdot[g]=[\xi(l(p_{v}))\triangleright g]\cdot[\xi(l(p_{v}))]^{-1}$ is satisfied for each $g\in c$. Hence, (4.9) holds by taking $g=(\xi(l(p_{v}))\triangleright)^{-1}l(p_{r})$. Finally, (4.10) is equivalent to property $(2)$ in the definition of rack § 4.1. ∎
@@
[…]
We wish to show the left vertical map is an equivalence, and so we will show the other three maps in the diagram are equivalences. The upper horizontal map is an equivalence by § 4.5, since the subrack is self-normalizing by § 4.5. The lower horizontal map is an equivalence because any homology class in $A^{\prime}[\alpha_{c^{\prime}}^{-1}]$ supported on a component that does not come from $\operatorname{Conf}^{c^{\prime}}$ is divisible by $\alpha_{c-c^{\prime}}$, and the completion is projection onto the summand where $\alpha_{c-c^{\prime}}$ acts by 00 by § 3.3. Finally, the right vertical map is an equivalence by § 4.5, where $c$ is replaced with $c^{\prime}$. ∎
@@ 5.2. A preliminary lemma on point counting
    Fix $i\geq 0$. There is some prime $\ell$, possibly depending on $n$, and a fixed $\mathbb{Q}_{\ell}$ vector space $V_{i}$ with the action of geometric Frobenius, $\operatorname{Frob}_{q}$, such that $V_{i}$ with its Frobenius action is independent of $n$. Assume there are constants $I$ and $J$, independent of $n$, so that whenever $n>Ii+J$, there is a fixed vector space $V_{i}$ with Frobenius action, so that $\operatorname{tr}(\operatorname{Frob}_{q}|V_{i})=\operatorname{tr}(\operatorname{Frob}_{q}|H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell}))$.
@@
$\left\|\frac{\left|Y_{n}(\mathbb{F}_{q})\right|}{q^{n}}-\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\right\|\leq\frac{2C^{\prime}}{1-\frac{C}{\sqrt{q}}}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}}.$
@@
$=$ $\sum_{0\leq i<\frac{n-J}{I}}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})+\sum_{i\geq\frac{n-J}{I}}(-1)^{i}\operatorname{tr}\left(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell}).\right)$
Subtracting $\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})$ from both sides of (5.1) and taking absolute values, we obtain
(5.2) $\left\|\frac{\left|Y_{n}(\mathbb{F}_{q})\right|}{q^{n}}-\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\right\|$
$\leq\sum_{i\geq\frac{n-J}{I}}\left\|\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\right\|+\sum_{i\geq\frac{n-J}{I}}(-1)^{i}\left\|\operatorname{tr}\left(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell})\right)\right\|$
We claim that $\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})$ converges absolutely as a function of $q$ for $\sqrt{q}>C$. Indeed, using Sun’s generalization of Deligne’s bounds for algebraic stacks [Sun12, Theorem 1.4] and our exponential bound assumption $(2)$ we obtain $\|\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\|\leq\frac{C^{\prime}C^{i}}{q^{i/2}}$. It follows that
$\sum_{i=0}^{\infty}\|(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\|\leq C^{\prime}\sum_{i=0}^{\infty}\left(\frac{C}{\sqrt{q}}\right)^{i}=C^{\prime}\frac{1}{1-\frac{C}{\sqrt{q}}}.$
@@
(5.3) $\sum_{i=\frac{n-J}{I}}^{\infty}\|(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\|\leq C^{\prime}\sum_{i=\frac{n-J}{I}}^{\infty}\left(\frac{C}{\sqrt{q}}\right)^{i}=C^{\prime}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}}\frac{1}{1-\frac{C}{\sqrt{q}}}.$
This bounds the first term on the right hand side of (5.2). It remains to bound the second term. Again, using Sun’s generalizations of Deligne’s bounds and our exponential bound assumption $(2)$, we similarly obtain $\|\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell}))\|\leq\frac{C^{\prime}C^{2n-j}}{q^{i/2}}$. Hence, so long as $C<\sqrt{q}$, we can bound
$\sum_{i\geq\frac{n-J}{I}}(-1)^{i}\left\|\operatorname{tr}\left(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell})\right)\right\|$ $\leq C^{\prime}\sum_{i\geq\frac{n-J}{I}}\left(\frac{C}{\sqrt{q}}\right)^{i}$

### v3

@@ 4.5. Showing the map is an equivalence
The proof we give here is modeled off of the argument in [RW20, Corollary 5.4]. We note that the argument given there has a small gap, where it is claimed on the first line of [RW20, p. 15] that $B\operatorname{Hur}^{c}_{G}\to B\operatorname{Conf}$ being a rational homology equivalence implies $\Omega B\operatorname{Hur}^{c}_{G}\to\Omega B\operatorname{Conf}$ is a rational homology equivalence on each path component. In general this does not hold, but it does hold when the spaces are weakly simple (see § 4.5 below), as was very helpfully pointed out to us by Oscar Randal-Williams. This observation allows us to fix the small gap.
@@
This lemma follows as in the proof of [RW20, Corollary 5.4]. Using § A.4 for $X=c$, $M=N=*_{+}$, we obtain an explicit topological model for $B\operatorname{Hur}^{c}$. By filtering this space by the number of labels on the interior, we get a filtration such that the associated cohomology spectral sequence is the cochain complex computing rack cohomology of $c$, see [EG03]. It follows from [EG03, Theorem 4.2] that the map $c\to*$ induces an isomorphism on rack cohomology with $\mathbb{Z}[\frac{1}{G^{0}_{c}}]$-coefficients, which completes the proof. ∎
Proof of § 4.5.
@@
We next wish to reduce to showing $B\operatorname{Hur}^{c}\to B\operatorname{Conf}^{c}$ is a rational homology equivalence. It is shown in [FRS07, Proposition 5.2] that for any rack $X$, BXBX is weakly simple. By § 4.5, $B\operatorname{Conf}^{c}$ is also weakly simple, since the action of $\pi_{1}$ on higher homotopy groups factors through $\pi_{1}B\operatorname{Conf}$ acting on the higher homotopy groups of $B\operatorname{Conf}$, and $B\operatorname{Conf}$ is the classifying space of a rack.
Applying § 4.5 to the map $B\operatorname{Hur}^{c}\to B\operatorname{Conf}^{c}$, we learn that if it is a a rational homology equivalence, then $\Omega B\operatorname{Hur}^{c}\to\Omega B\operatorname{Conf}^{c}$ is also a rational homology equivalence. Here, we are also using that the map $\pi_{1}B\operatorname{Hur}^{c}\to\pi_{1}B\operatorname{Conf}^{c}$ is an equivalence by construction of $\operatorname{Conf}^{c}$.
We next claim that $B\operatorname{Conf}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology. The map $\Omega B\pi_{0}\operatorname{Hur}^{c}\to\Omega B\pi_{0}\operatorname{Conf}$ is surjective with kernel a group we call $G$. Here, $G$ is finite because of the fact that $|\pi_{0}\operatorname{Conf}^{c}_{n}|=|\pi_{0}\operatorname{Hur}^{c}_{n}|$ stabilizes as $n$ grows, which is a consequence of § 4.4 and the fact that a filtered colimit of finite sets along surjections is finite. It follows from this and the fact that $\psi$ from (4.3) is an equivalence that the fiber of the map $B\operatorname{Conf}^{c}\to B\operatorname{Conf}$ is $BG$. We then obtain our claim that $B\operatorname{Conf}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology using the Serre spectral sequence for this fiber sequence, since the rational homology of $G$ is just $\mathbb{Q}$ in degree 0 with trivial action of $B\operatorname{Conf}$.
It remains then to show that the map $B\operatorname{Hur}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology, but this follows from applying § 4.5, since c/c=c/c=* since $c$ is connected, and $\operatorname{Hur}^{*}$ is $\operatorname{Conf}$. ∎
@@
For $c^{\prime}\subset c$ a subrack, there is a map of $\mathbb{E}_{1}$-algebras $\overline{i}_{c^{\prime}}^{c}:\operatorname{Conf}^{c^{\prime}}\to\operatorname{Conf}^{c}$ that is an inclusion of components. Similarly, there is a map $\tilde{i}_{c^{\prime}}^{c}:\operatorname{Hur}^{c^{\prime}}\to\operatorname{Hur}^{c}$. We have a section of the induced map $\operatorname{Conf}^{c^{\prime}}_{+}\to\operatorname{Conf}^{c}_{+}$ of pointed $\mathbb{E}_{1}$-algebras $\overline{r}_{c}^{c^{\prime}}:\operatorname{Conf}^{c}_{+}\to\operatorname{Conf}^{c^{\prime}}_{+}$ obtained by sending all components not in image of $\operatorname{Conf}^{c}$ to the base point. Similarly there is a map $\widetilde{r}_{c}^{c^{\prime}}:\operatorname{Hur}^{c}_{+}\to\operatorname{Hur}^{c^{\prime}}_{+}$ obtained the same way. Let $r_{c}^{c^{\prime}}$ denote the associated $\mathbb{E}_{1}$-algebra map $A\to C_{*}(\operatorname{Hur}^{c^{\prime}})$ obtained by taking reduced chains. In particular, this map of algebras sends all generators $\alpha_{x}$, for $x\in c-c^{\prime}$, to 0.
@@
We first note that since $\alpha_{c-c^{\prime}}$ acts by 0 on $C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$, completion doesn’t change it, so the map $A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}\to C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$ in the proposition statement is obtained by taking $r_{c}^{c^{\prime}}$, localizing at $\alpha_{c^{\prime}}^{-1}]$ and completing at $\alpha_{c-c^{\prime}}$.
@@
[…]
@@
Note that all of the objects in the triangle are connective. By § 3.3, $\pi_{0}A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ is the quotient of $\pi_{0}A[\alpha_{c^{\prime}}^{-1}]^{\wedge}$ by the two sided ideal generated by $\alpha_{c-c^{\prime}}$. This ideal is exactly the classes coming from components not in $\operatorname{Hur}^{c^{\prime}}$. Thus the map $h$ is an equivalence on $\pi_{0}$. Moreover, $A\to A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ is a hom. epi. by § 4.4 and § 3.3.
@@
is an ind-weak homology equivalence (see § A.4) as $\epsilon$ approaches 0 where $0<\epsilon<1$.
We now recall the description of these spaces, which follows from § A.4. First, ${Q}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$ is the space whose points consist of tuples $(m,n,x,l:x\to c)$ with the following properties: $m\in X_{+},n\in X_{+}$, $x$ is a finite subset $x$ of $0,1]\times[\epsilon,1-\epsilon]$ such that if $\pi_{2}$ is the projection onto the second coordinate, then $\pi_{2}$ is injective on $x$, and the points in $\pi_{2}(x)$ are distance at least $\epsilon$ from each other, and $l:x\to c$ is a map of sets, which we colloquially refer to as a labeling function.
@@
Suppose $y$ as above lies in $F_{j}Q_{\epsilon}^{c,c^{\prime}}$. Define $v$ to be the largest number so that $l(p_{v})\notin N_{c}(c^{\prime})$. By definition of the filtration $F_{j}Q_{\epsilon}^{c,c^{\prime}}$, we have $v\geq k-j$. Let $\epsilon^{\prime}:=\frac{\epsilon}{2}$. We now construct a continuous homotopy $H:F_{j}Q^{c,c^{\prime}}_{\epsilon}\times I\to G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$ as follows: If either $m,n$ is the base point, the point $y$ is equivalent to the base point under relation $(iii)$, and we choose the constant homotopy at the base point. So now assume this is not the case. If v>kjv>k-j, then the point is in $F_{j-1}Q_{\epsilon}^{c,c^{\prime}}$, so goes to the base point in the associated graded $G_{j}Q_{\epsilon}^{c,c^{\prime}}$. Thus, we may choose the constant homotopy on such points as well. So now we may assume as well that v=kjv=k-j.
@@
It only remains to consider then the case that the point $p_{r}$ being removed in $(i)$ has $l(p_{r})\in c^{\prime}$. Note that we may assume $r\neq v$ because $l(p_{v})\notin c^{\prime}$. Because we are working on the subquotient $G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$, if we apply $(i)$ for some point $p_{r}$ with $r>v$, the point becomes identified with the base point, and so $H$ is the constant homotopy on the base point. (This is the reason for introducing the filtration $F_{j}$.) Hence, it only remains to deal with the case that $r<v$ at a point $y$ presented as $(m,n,x,l:x\to c)$, as in $(\dagger)$ above.
@@
First, recall the defining relation of Hurwitz space from § 4.1 implies that the positive generator of the braid group identifies $g]\cdot[h]=[h]\cdot[h\triangleright g]$. Therefore, taking $h=\xi(l(p_{v}))$, and using that $\xi(l(p_{v}))]$ acts invertibly on $M$, the action on $M$ factors through a group where the relation $\xi(l(p_{v}))]^{-1}\cdot[g]=[\xi(l(p_{v}))\triangleright g]\cdot[\xi(l(p_{v}))]^{-1}$ is satisfied for each $g\in c$. Hence, (4.9) holds by taking $g=(\xi(l(p_{v}))\triangleright)^{-1}l(p_{r})$. Finally, (4.10) is equivalent to property $(2)$ in the definition of rack § 4.1. ∎
@@
$A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$$C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$$A^{\prime}[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$$C_{*}(\operatorname{Conf}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}].$}
We wish to show the left vertical map is an equivalence, and so we will show the other three maps in the diagram are equivalences. The upper horizontal map is an equivalence by § 4.5, since the subrack is self-normalizing by § 4.5. The lower horizontal map is an equivalence because any homology class in $A^{\prime}[\alpha_{c^{\prime}}^{-1}]$ supported on a component that does not come from $\operatorname{Conf}^{c^{\prime}}$ is divisible by $\alpha_{c-c^{\prime}}$, and the completion is projection onto the summand where $\alpha_{c-c^{\prime}}$ acts by 0 by § 3.3. Finally, the right vertical map is an equivalence by § 4.5, where $c$ is replaced with $c^{\prime}$. ∎
@@ 5.2. A preliminary lemma on point counting
    There are real numbers $I$ and $J$, only depending on the sequence $\{Y_{n}\}_{n\in S}$ with the following property. Suppose there is some fixed $\ell$ and a collection of complex numbers $t_{i}$ so that whenever $n>Ii+J$, $\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell}))=t_{i}$.
@@
$\left\|\frac{\left|Y_{n}(\mathbb{F}_{q})\right|}{q^{n}}-\sum_{i=0}^{\infty}(-1)^{i}t_{i}\right\|\leq\frac{2C^{\prime}}{1-\frac{C}{\sqrt{q}}}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}}.$
@@
$=$ $\sum_{0\leq i<\frac{n-J}{I}}(-1)^{i}t_{i}+\sum_{i\geq\frac{n-J}{I}}(-1)^{i}\operatorname{tr}\left(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell}).\right)$
Subtracting $\sum_{i=0}^{\infty}(-1)^{i}t_{i}$ from both sides of (5.1) and taking absolute values, we obtain
(5.2) $\left\|\frac{\left|Y_{n}(\mathbb{F}_{q})\right|}{q^{n}}-\sum_{i=0}^{\infty}(-1)^{i}t_{i}\right\|$
[…]
We claim that $\sum_{i=0}^{\infty}(-1)^{i}t_{i}$ converges absolutely as a function of $q$ for $\sqrt{q}>C$. Indeed, using Sun’s generalization of Deligne’s bounds for algebraic stacks [Sun12, Theorem 1.4] and our exponential bound assumption $(2)$ we obtain $t_{i}\leq\frac{C^{\prime}C^{i}}{q^{i/2}}$. It follows that
$\sum_{i=0}^{\infty}\|(-1)^{i}t_{i}\|\leq C^{\prime}\sum_{i=0}^{\infty}\left(\frac{C}{\sqrt{q}}\right)^{i}=C^{\prime}\frac{1}{1-\frac{C}{\sqrt{q}}}.$
@@
(5.3) $\sum_{i=\frac{n-J}{I}}^{\infty}\|(-1)^{i}t_{i}\|\leq C^{\prime}\sum_{i=\frac{n-J}{I}}^{\infty}\left(\frac{C}{\sqrt{q}}\right)^{i}=C^{\prime}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}}\frac{1}{1-\frac{C}{\sqrt{q}}}.$
[…]